OpenStudy (anonymous):

Can anyone help me with this graph problem? I have to use either the elimination or substitution to find the solutions (if any). And I have to graph the two lines, and list the x-intercepts, y-intercepts, and points of intersection. Here is the first one.......... 3x + 7y = 3 and 12x + 28y = 16

OpenStudy (anonymous):

Okay, so first, solve both equations for y: y = (3 - 3x)/7 y = (16 - 12x)/28 Next, set y = y: (3 - 3x)/7 =(16 - 12x)/28 Next, solve for x by cross multiplying and grouping methods: 28(3-3x) = 7(16-12x) 84 - 84x = 112 - 84x Hmmm, I get no solution at this step...

OpenStudy (anonymous):

There is no solution, therefore, there is no need to do anything further

OpenStudy (anonymous):

ok, how would you show that on a graph and how do I determine the x-intercept, y-intercept and slope?

OpenStudy (anonymous):

Q: How do you show that on a graph? A: Graph the following: y = (3 - 3x)/7 y = (16 - 12x)/28 Q: How do you determine the x-intercept? A: To determine the x-intercept, set y = 0, then solve for x Q: How do you determine the y-intercept? A: Set x = 0, solve for y Q: How do you find the slope? A: Observe The following Equations: y = (3 - 3x)/7 y = (16 - 12x)/28 Notice that the slope for the first one is 3x/7 Notice that the slope for the second one is 12x/28 or 3x/7 So the slope for both is 3x/7

OpenStudy (anonymous):

If an equation has the same slope, the graph of both are parallel, thus no intersection, and thus no solution

OpenStudy (anonymous):

Thank you for trying Bliss, but I have no idea what I am doing on this

OpenStudy (anonymous):

I'm sorry to hear that....

OpenStudy (anonymous):

The only thing I could try is post an online classroom with voice and explain it to you. Are you interested?